There are two carotid arteries that feed blood to
Answer #1

Rate flow = pressure drop / flow resistance

(Flow resistance)left=  \frac{8\eta(9.2x10^{-2})}{\pi(2.65x10^{-3})^4}+\frac{8\eta(2x10^{-2})}{\pi(1.6x10^{-3})^4}

(Flow resistance)right=  \frac{8\eta(11.2x10^{-2})}{\pi(2.65x10^{-3})^4}

For the same flow rate

\frac{\Delta P_{left}}{\Delta P_{right}}=\frac{ \frac{9.2}{(2.65)^4}+ \frac{2}{(1.6)^4}}{\frac{11.2}{(2.65)^4}}

\frac{\Delta P_{left}}{\Delta P_{right}}=2.17